# On the Sharp Input-Output Analysis of Nonlinear Systems under Adversarial Attacks

### v1

@@ Appendix B Preliminaries on sub-Gaussian variables
For $i\in\{1,\dots,r\}$, define $\beta_{i}(\Delta_{i}):=\sum_{t=\tau}^{T-1}|\Delta_{i}^{T}\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau,i)}|-|\bm{W}_{t}^{(\tau,i)}|$. Recall from (D), (38), and (58) that whenever the norm of $\Delta_{i}$ is sufficiently small, we have
$\beta_{i}(\Delta_{i})=\sum_{t=\tau}^{T-1}z_{t}^{i}(\Delta_{i})=\|\Delta_{i}\|_{2}\sum_{t=\tau}^{T-1}z_{t}^{i}\Bigr{(}\frac{\Delta_{i}}{\|\Delta_{i}\|_{2}}\Bigr{)},$ (86)
where $z_{t}^{i}(\cdot)$ from (58) satisfies absolute homogeneity. Due to the convexity of $\beta_{i}(\cdot)$, for all $h\geq 1$ and $s,\tilde{s}\in\mathbb{S}^{M-1}$, we have
$\beta_{i}\Bigr{(}\frac{1}{h}s+(1-\frac{1}{h})\tilde{s}\Bigr{)}\leq\frac{1}{h}\beta_{i}(s)+\Bigr{(}1-\frac{1}{h}\Bigr{)}\beta_{i}(\tilde{s}).$
Substituting $s=h\Delta_{i}$ and $\tilde{s}=0$ incurs
$\beta_{i}(h\Delta_{i})\geq h\beta_{i}(\Delta_{i})=\|h\Delta_{i}\|_{2}\sum_{t=\tau}^{T-1}z_{t}^{i}\Bigr{(}\frac{\Delta_{i}}{\|\Delta_{i}\|_{2}}\Bigr{)}$ (87)
when $\Delta_{i}$ is sufficiently small. Since the inequality (87) holds for all $h\geq 1$, it implies that we generally have
$\beta_{i}(\Delta_{i})\geq\|\Delta_{i}\|_{2}\sum_{t=\tau}^{T-1}z_{t}^{i}\Bigr{(}\frac{\Delta_{i}}{\|\Delta_{i}\|_{2}}\Bigr{)}\geq\|\Delta_{i}\|_{2}\cdot\Theta\left(\frac{(2(1-p)^{\tau}-1)\lambda T}{\tau\nu^{3}}\right)$ (88)
with probability at least $1-\frac{\delta}{2}$, regardless of the magnitude of $\Delta_{i}$. The last inequality comes from $\frac{\Delta_{i}}{\|\Delta_{i}\|_{2}}\in\mathbb{S}^{M-1}$, in which we can apply Corollary D.10 given the time (84).
In the meantime, the optimality of $\hat{G}_{T}$ to the $\ell_{1}$-norm estimator (12) for the system (11) yields
$\hat{G}_{T}=\operatorname*{arg\,min}_{G}\sum_{t=\tau}^{T-1}\left\|(G^{*}-G)\cdot\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}\right\|_{1},$
@@
$\sum_{t=\tau}^{T-1}\|(G^{*}-\hat{G}_{T})\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}\|_{1}-\|\bm{x}_{t}^{(\tau)}\|_{1}$ (89)
$\leq\sum_{t=\tau}^{T-1}\|(G^{*}-\hat{G}_{T})\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}\|_{1}\leq\sum_{t=\tau}^{T-1}\|\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}\|_{1}$ (90)
$\leq\sum_{t=\tau}^{T-1}\|\bm{W}_{t}^{(\tau)}\|_{1}+\|\bm{x}_{t}^{(\tau)}\|_{1},$ (91)
where (90) uses the optimality of $\hat{G}_{T}$ and the other inequalities are from the triangle inequality.
For $i\in\{1,\dots,r\}$, let $g_{i}^{*}$ and $\hat{g}_{i}$ denote the $i^{\text{th}}$ rows of $G^{*}$ and $\hat{G}_{T}$, respectively. Then, we have
$\sum_{i=1}^{r}\beta_{i}(g_{i}^{*}-\hat{g}_{i})=\sum_{t=\tau}^{T-1}\|(G^{*}-\hat{G}_{T})\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}\|_{1}-\|\bm{W}_{t}^{(\tau)}\|_{1}\leq 2\sum_{t=\tau}^{T-1}\|\bm{x}_{t}^{(\tau)}\|_{1},$ (92)
where the equality is from the definition of $f_{i}(\cdot)$ and the inequality is by (89) and (91). Recall from Lemma 2.5 that $\|\bm{x}_{t}^{(\tau)}\|_{2}\leq CL\rho^{\tau}\|\bm{x}_{t-\tau}\|_{2}$. Then, we can establish that
$2\sum_{t=\tau}^{T-1}\|\bm{x}_{t}^{(\tau)}\|_{1}\leq 2\sqrt{r}\sum_{t=\tau}^{T-1}\|\bm{x}_{t}^{(\tau)}\|_{2}\leq 2\sqrt{r}\sum_{t=\tau}^{T-1}CL\rho^{\tau}\|\bm{x}_{t-\tau}\|_{2}=2\sqrt{r}\sum_{t=0}^{T-\tau-1}CL\rho^{\tau}\|\bm{x}_{t}\|_{2}.$ (93)
Thus, we have
$\sum_{i=1}^{r}\beta_{i}(g_{i}^{*}-\hat{g}_{i})\leq 2\sqrt{r}\sum_{t=0}^{T-\tau-1}CL\rho^{\tau}\|\bm{x}_{t}\|_{2}.$
Given the time (84), we are able to lower bound the left-hand side with probability at least $1-\frac{\delta}{2}$ using (88), and upper bound the right-hand side with probability at least $1-\frac{\delta}{2}$ using Lemma D.11. This can be written as
$\sum_{i=1}^{r}\|g_{i}^{*}-$ $\hat{g}_{i}\|_{2}\cdot\Theta\left(\frac{(2(1-p)^{\tau}-1)\lambda T}{\tau\nu^{3}}\right)\leq 2\sqrt{r}CL\rho^{\tau}\cdot\Theta\left(\frac{(\sigma_{u}+\sigma_{w})}{1-\rho}\cdot T\right),$ (94)
@@
$\sum_{i=1}^{r}\|g_{i}^{*}-$ $\hat{g}_{i}\|_{2}\leq\Theta\left(\frac{\rho^{\tau}L(\sigma_{u}+\sigma_{w})\tau\nu^{3}\sqrt{r}}{(1-\rho)(2(1-p)^{\tau}-1)\lambda}\right).$
The relationship $\|G^{*}-\hat{G}_{T}\|_{F}\leq\sum_{i=1}^{r}\|g_{i}^{*}-\hat{g}_{i}\|_{2}$ completes the proof. ∎
@@
Since we have $-\log(1-x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots\leq x+x^{2}+x^{3}+\cdots=\frac{x}{1-x}<2x$ for $|x|<\frac{1}{2}$, it follows that
$l\geq\frac{\log\left(T/\delta\right)}{\frac{2}{2\tau+1}}=\Theta\left(\tau\log\left(\frac{T}{\delta}\right)\right).$
@@
$\mathbb{P}\left(M_{T}<\Theta\left(\tau\log\left(\frac{T}{\delta}\right)\right)\right)\geq 1-\delta$ (96)
@@
Suppose that control inputs $u_{t}$ are chosen independently from $\{-1,1\}$ with equal probability for all $t=0,\dots,T-1$, which satisfies Assumption 2.3. Given a finite $\sigma_{w}\geq\bigr{(}\frac{1}{\rho}\bigr{)}^{\Theta(\tau\log(T/\delta))}$, start the system with $x_{0}=\sigma_{w}$ and let the disturbance $w_{t}$ also be $\sigma_{w}$ whenever the attack occurs at each time $t$, which satisfies Assumption 2.4. Note that the dynamics $f$ shrinks the system by a factor of $\rho$. Then, considering (96), one can ensure that adversarial attacks yield $x_{t}\geq 1$ for all $t=0,\dots,T-1$ with probability at least $1-\delta$. In this case, we can also rewrite (E) as:
@@
[…]
@@
where one has $10<c<$ 1 under $10<\rho<$ 1. Then, the absolute difference of (101a) and (101b) is calculated as
@@
since $u_{t-\tau}$ is selected from $-1$ and $1$, and $\beta(x)$ is an odd function. Now, let the basis function be
@@
which consists of (101a) and (101b). Since the expression (101a) is the input-output mapping of the true system (E), the true matrix $G^{*}$ in (11) is $[1~{}~{}0]$. However, we again recall that under adversarial attacks, any estimator cannot distinguish (99) from (E), and may instead recover the input-output mapping of the alternative system (99), resulting in the estimate $\hat{G}_{T}=[0~{}~{}1]$. This always leads to an estimation error of $\sqrt{2}$.
@@
$\frac{L\rho^{\tau}}{\lambda}\cdot\sqrt{\frac{2}{3}}c=\Theta\left(\frac{L\rho^{\tau}}{\lambda}\right).$
Note that this lower bound was obtained from a scalar system. Considering an $r$-dimensional $y_{t}$, where each coordinate follows the same scalar system as discussed, the lower bound on the estimation error becomes $\Theta\bigr{(}\frac{L\rho^{\tau}\sqrt{r}}{\lambda}\bigr{)}$, since the Frobenius norm is used to measure the error. ∎
@@ Appendix F Numerical Experiment Details
In this section, we will present experiment details on Section 4. Apple M1 Chip with 8-Core CPU is sufficient for the experiments. The error bars (shaded area) in all the figures in the paper report 95% confidence intervals based on the standard error. We calculate the standard error by running 10 different experiments by generating 10 random sets of matrices $A,B,C$, and $D$ and using random adversarial disturbances for each experiment.
@@
Both functions are symmetric around the origin, monotonic, and $1$-Lipschitz, which are desirable for activation functions of a neural net. Note that the first function is bounded within $[-1,1]$, while the second function is unbounded. We analyze both options to determine whether the boundedness affects the behavior of the estimation error.
@@
Our first experiment compares the $\ell_{1}$-norm estimator with the commonly used least-squares under $\tau=5$ and $\rho=0.5$. We consider both cases of control inputs given in (105). Based on the observation trajectory, we evaluate the following two estimators using the MOSEK solver [MOSEK ApS, 2025]:
$\operatorname*{arg\,min}_{G}\sum_{t=\tau}^{T-1}\left\|y_{t}-G\cdot\Phi(\bm{U}_{t}^{(\tau)})\right\|_{1}\quad\text{vs.}\quad\operatorname*{arg\,min}_{G}\sum_{t=\tau}^{T-1}\left\|y_{t}-G\cdot\Phi(\bm{U}_{t}^{(\tau)})\right\|_{2}^{2}.$
Our second experiment analyzes the effect of $\tau$ and $\rho$ on the $\ell_{1}$-norm estimator under zero-mean Gaussian inputs. The third experiment analyzes the same effect under nonzero-mean uniform inputs. For both experiments, we consider $\tau\in[5,10]$ and $\rho\in[0.25,0.5]$. Note that all experiments were conducted for both activation functions given in (103).
(a) Comparison of Least-squares and $\ell_{1}$-norm estimator
(b) Analysis of $\tau$ and $\rho$ effects under Gaussian inputs
@@
The experiments using $\sigma(x)=\tanh(x)$ are shown in Figure 1 and those with $\sigma(x)=\text{sgn}(x)\cdot\log(|x|+1)$ are presented in Figure 2. As noted earlier, the two functions differ in their boundedness. In both Figures 1(a) and 2(a), one can observe that the $\ell_{1}$-norm estimator accommodates both Gaussian and uniform inputs and arrive at a similar stable region, unlike the least-squares estimator. The discrepancy in the estimation error with respect to $\tau$ and $\rho$ can clearly be observed in Figures 1(b), 1(c), 2(b), and 2(c), where the estimation error decreases as $\tau$ increases and $\rho$ decreases, which is consistent with our theoretical optimal error of $\Theta(\rho^{\tau})$. These findings remain valid regardless of the boundedness of the activation function $\sigma$.

### v2

@@ Appendix B Preliminaries on sub-Gaussian variables
The optimality of $\hat{G}_{T}$ to the $\ell_{2}$-norm estimator (12) for the system (8) yields
$\hat{G}_{T}=\operatorname*{arg\,min}_{G}\sum_{t=\tau}^{T-1}\left\|(G^{*}-G)\cdot\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}+\bm{\epsilon}_{t}\right\|_{2},$
@@
$\sum_{t=\tau}^{T-1}\|(G^{*}-\hat{G}_{T})\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}\|_{2}-\|\bm{x}_{t}^{(\tau)}+\bm{\epsilon}_{t}\|_{2}$ (69)
$\leq\sum_{t=\tau}^{T-1}\|(G^{*}-\hat{G}_{T})\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}+\bm{\epsilon}_{t}\|_{2}\leq\sum_{t=\tau}^{T-1}\|\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}+\bm{\epsilon}_{t}\|_{2}$ (70)
$\leq\sum_{t=\tau}^{T-1}\|\bm{W}_{t}^{(\tau)}\|_{2}+\|\bm{x}_{t}^{(\tau)}+\bm{\epsilon}_{t}\|_{2},$ (71)
where (70) uses the optimality of $\hat{G}_{T}$ and the other inequalities are from the triangle inequality. By rearranging, we have
$\sum_{t=\tau}^{T-1}\|(G^{*}-\hat{G}_{T})\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}\|_{2}-\|\bm{W}_{t}^{(\tau)}\|_{2}\leq 2\sum_{t=\tau}^{T-1}\|\bm{x}_{t}^{(\tau)}\|_{2}+\|\bm{\epsilon}_{t}\|_{2},$ (72)
where the inequality is by (69) and (71). Recall from Lemma 2.3 that $\|\bm{x}_{t}^{(\tau)}\|_{2}\leq CL\rho^{\tau}\|\bm{x}_{t-\tau}\|_{2}$. Then, we can establish that
$2\sum_{t=\tau}^{T-1}\|\bm{x}_{t}^{(\tau)}\|_{2}\leq 2\sum_{t=\tau}^{T-1}CL\rho^{\tau}\|\bm{x}_{t-\tau}\|_{2}=2\sum_{t=0}^{T-\tau-1}CL\rho^{\tau}\|\bm{x}_{t}\|_{2}.$ (73)
Given the time (67), the right-hand side of (72) is upper bounded by
$2\cdot O\left(\left(\frac{CL\rho^{\tau}(\sigma_{u}+\sigma_{w})}{1-\rho}+\bar{\epsilon}\right)T\right)$
with probability at least $1-\frac{\delta}{2}$, which follows from Lemma D.9 and $\|\bm{\epsilon}_{t}\|_{2}\leq\bar{\epsilon}$.
We now aim to lower bound the left-hand side of (72) given the time (67). Inspired by (33), we have
$\sum_{t=\tau}^{T-1}\|(G^{*}-\hat{G}_{T})\Phi(\bm{U}_{t}^{(\tau)})$ $+\bm{W}_{t}^{(\tau)}\|_{2}-\|\bm{W}_{t}^{(\tau)}\|_{2}\geq\sum_{t=\tau}^{T-1}\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|(G^{*}-\hat{G}_{T})\cdot\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}$
$=\|G^{*}-\hat{G}_{T}\|_{F}\cdot\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\Biggr\|\frac{G^{*}-\hat{G}_{T}}{\|G^{*}-\hat{G}_{T}\|_{F}}\cdot\Phi(\bm{U}_{t}^{(\tau)})\Biggr\|_{2}$
$=\|G^{*}-\hat{G}_{T}\|_{F}\cdot\Omega\left(\frac{(2(1-p)^{\tau}-1)\cdot\lambda T}{4\nu^{3}}\right)$
where the first equality comes from the homogeneity of the $\ell_{2}$-norm, and the second equality holds for any $G^{*}-\hat{G}_{T}$ with probability at least $1-\frac{\delta}{2}$ due to Theorem D.8.
Thus, with probability at least $1-\delta$, we have
$\|G^{*}-\hat{G}_{T}\|_{F}\cdot\Omega\left(\frac{(2(1-p)^{\tau}-1)\cdot\lambda T}{4\nu^{3}}\right)\leq 2\cdot O\left(\left(\frac{CL\rho^{\tau}(\sigma_{u}+\sigma_{w})}{1-\rho}+\bar{\epsilon}\right)T\right),$
@@
$\|G^{*}-\hat{G}_{T}\|_{F}=O\left(\left(\frac{\rho^{\tau}L(\sigma_{u}+\sigma_{w})}{1-\rho}+\bar{\epsilon}\right)\cdot\frac{\nu^{3}}{(2(1-p)^{\tau}-1)\lambda}\right).$
This completes the proof. ∎
@@
Since we have $\log(1-x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots\leq x+x^{2}+x^{3}+\cdots=\frac{x}{1-x}<2x$ for $|x|<\frac{1}{2}$, it follows that
$l\geq\frac{\log\left(T/\delta\right)}{\frac{2}{2\tau+1}}\geq\tau\log\left(\frac{T}{\delta}\right).$
@@
$\mathbb{P}\left(M_{T}<\tau\log\left(\frac{T}{\delta}\right)\right)\geq 1-\delta$ (75)
@@
Suppose that control inputs $u_{t}$ are chosen independently from $\{-1,1\}$ with equal probability for all $t=0,\dots,T-1$, which satisfies Assumption 2.5. Given a finite $\sigma_{w}=\bigr(\frac{1}{\rho}\bigr)^{\Omega(\tau\log(T/\delta))}$, start the system with $x_{0}=\sigma_{w}$ and let the disturbance $w_{t}$ also be $\sigma_{w}$ whenever the attack occurs at each time $t$, which satisfies Assumption 2.6. Note that the dynamics $f$ shrinks the system by a factor of $\rho$. Then, considering (75), one can ensure that adversarial attacks yield $x_{t}\geq 1$ for all $t=0,\dots,T-1$ with probability at least $1-\delta$. In this case, we can also rewrite (E) as:
@@
$\beta(x)=\begin{dcases*}\frac{\tanh(x)}{\tanh(1)},&if $-1\leq x\leq 1$,\\ x,&otherwise,\end{dcases*}$ (79)
@@
where one has $0<c<1$ under $0<\rho<1$. Then, the absolute difference of (80a) and (80b) is calculated as
@@
since $u_{t-\tau}$ is selected from 1-1 and $1$, and $\beta(x)$ is an odd function. Now, let the basis function be
@@
which consists of (80a) and (80b). This implies that the approximation error vector is designed to be $\bm{\epsilon}_{t}=0$.
Since the expression (80a) is the input-output mapping of the true system (E), the true matrix $G^{*}$ in (8) is [10][1~~0]. However, we again recall that under adversarial attacks, any estimator cannot distinguish (78) from (E), and may instead recover the input-output mapping of the alternative system (78), resulting in the estimate $\hat{G}_{T}=[0~~1]$. This always leads to an estimation error of $\sqrt{2}$.
@@
$\frac{L\rho^{\tau}}{\lambda}\cdot\sqrt{\frac{2}{3}}c=\Omega\left(\frac{L\rho^{\tau}}{\lambda}\right),$
which completes the proof. ∎
@@ Appendix F Numerical Experiment Details
In this section, we will present experiment details on Section 4. Apple M1 Chip with 8‑Core CPU is sufficient for the experiments. The error bars (shaded area) in all the figures in the paper report 95% confidence intervals based on the standard error. We calculate the standard error by running 10 different experiments by generating 10 random sets of matrices $A,B,C$, and $D$ and using random adversarial disturbances for each experiment.
@@
Both functions are symmetric around the origin, monotonic, and $1$-Lipschitz, which are desirable for activation functions of a neural net. Note that the first function is bounded within [1,1][-1,1], while the second function is unbounded. We analyze both options to determine whether the boundedness affects the behavior of the estimation error.
@@
Our first experiment compares the $\ell_{2}$-norm estimator with the commonly used least-squares under $\tau=5$ and $\rho=0.5$. We consider both cases of control inputs given in (84). Based on the observation trajectory, we evaluate the following two estimators using the MOSEK solver (MOSEK ApS, 2025):
$\operatorname*{arg\,min}_{G}\sum_{t=\tau}^{T-1}\left\|y_{t}-G\cdot\Phi(\bm{U}_{t}^{(\tau)})\right\|_{2}\quad\text{vs.}\quad\operatorname*{arg\,min}_{G}\sum_{t=\tau}^{T-1}\left\|y_{t}-G\cdot\Phi(\bm{U}_{t}^{(\tau)})\right\|_{2}^{2}.$
Our second experiment additionally compares the $\ell_{2}$-norm estimator with the $\ell_{1}$-norm estimator and the $\ell_{\infty}$-norm estimator:
$\operatorname*{arg\,min}_{G}\sum_{t=\tau}^{T-1}\left\|y_{t}-G\cdot\Phi(\bm{U}_{t}^{(\tau)})\right\|_{1}\quad\text{and}\quad\operatorname*{arg\,min}_{G}\sum_{t=\tau}^{T-1}\left\|y_{t}-G\cdot\Phi(\bm{U}_{t}^{(\tau)})\right\|_{\infty}.$
Our third experiment analyzes the effect of $\tau$ and $\rho$ on the $\ell_{2}$-norm estimator under nonzero-mean uniform inputs, where we consider $\tau\in[5,10]$ and $\rho\in[0.25,0.5]$. Note that all experiments were conducted for both activation functions given in (82).
(a) Comparison of Least-squares and $\ell_{2}$-norm estimator
(b) Comparison of the $\ell_{\alpha}$-norm estimators ($\alpha=1,2,\infty$)
@@
The experiments using $\sigma(x)=\tanh(x)$ are shown in Figure 2 and those with $\sigma(x)=\text{sgn}(x)\cdot\log(|x|+1)$ are presented in Figure 3. As noted earlier, the two functions differ in their boundedness. In both Figures 2(a) and 3(a), one can observe that the $\ell_{2}$-norm estimator accommodates both Gaussian and uniform inputs and arrive at a similar stable region, unlike the least-squares estimator.
Furthermore, both Figures 2(b) and 3(b) demonstrate that all norm estimators accurately recover the true matrix $G^{*}$, with the $\ell_{2}$-norm estimator achieving the smallest error among them. This supports the findings in Remark 3.4, which states that only the $\ell_{2}$-norm estimator attains the optimal error (14) that matches the lower bound presented in Theorem 3.5.
The discrepancy in the estimation error with respect to $\tau$ and $\rho$ can clearly be observed in Figures 2(c) and 3(c), where the estimation error using the $\ell_{2}$-norm estimator decreases as $\tau$ increases and $\rho$ decreases, which is consistent with our theoretical optimal error of $O(\rho^{\tau})$. These findings remain valid regardless of the boundedness of the activation function $\sigma$.